OpenStudy (nermaljean99):

A biologist took count of the number of fish in a particular lake and recounted the lake's population of fish each of the next six weeks. Week: 0, 1, 2, 3, 4, 5, 6 Population: 350, 353, 382, 437, 518, 625, 758 Find a quadratic function that models the data as a function of x, the number of weeks. Use the model to estimate the number of fish at the lake on week 8.

OpenStudy (imstuck):

Are you allowed to use the ExpReg function on your calculator to find this?

OpenStudy (confluxepic):

Yeah. Sometimes you are not allowed to use it.

OpenStudy (nermaljean99):

Uhh I dont know would it help if I put the answers?

OpenStudy (imstuck):

I got this as an equation using the ExpReg function on my calculator: \[y=313.9275644*(1.143939023)^{x}\]It looks completely crazy, but if you enter any number of weeks for x, you will get the corresponding y value. If I enter 8 for x, the y value I get, the population of fish, is 920.5

OpenStudy (nermaljean99):

A.) P(x)= 13x^2-10x+350; 917 fish B.) P(x)= 13x^2-10x+350; 1102 fish C.) P(x)= 18x^2+10x+300; 1252 fish D.) P(x)= 18x^2+10x+300; 1532 fish

OpenStudy (nermaljean99):

@IMStuck

OpenStudy (nermaljean99):

@KlOwNlOvE

OpenStudy (nermaljean99):

@hugsnotughs think you could help me again? :)

OpenStudy (hugsnotughs):

I'm acually about to leave, sorry. :/